27. Film critics gave the film Basic Statistics: The Movie an average rating of 8.1 (on a scale of 0-10). The standard deviation is 0.7. You drew a random sample of n = 56 from all film critics and asked them to rate the film Boyhood with a number. What is the probability that the average rating in the sample is greater than 8.0?
• 14%
• 44%
• 86%
• 56%

Answer by stanbon(75887)

Film critics gave the film Basic Statistics: The Movie an average rating of 8.1 (on a scale of 0-10). The standard deviation is 0.7. You drew a random sample of n = 56 from all film critics and asked them to rate the film Boyhood with a number. What is the probability that the average rating in the sample is greater than 8.0?
------
z(8) = (8-8.1)/[0.7/sqrt(56)] = -1.0690
P(x-bar > 8) = P(z > -1.0690) = normalcdf(-1.0690,100) = 86%
---------------
Cheers,
Stan H.
------------------
• 14%
• 44%
• 86%
• 56%